OpenStudy (anonymous):

2. The manager of a fish store has water that is 10% salt and water that is 25% salt. He needs to fill an aquarium with 5 gallons of water that is 20% salt. a. Write a system of linear equations that you can use to determine how many gallons of each type of salt water the manager should combine. Be sure to define your variables. x=10% of salt water and y=25% of salt water so x+y=5 b. Solve the system and determine how many gallons of each type of salt water the manager should combine. Show all your work.

OpenStudy (tkhunny):

"x=10% of salt water and y=25% of salt water so x+y=5" Not bad. Let me suggest more clear defintiions. x = The amount of 10% salt water to be added to the tank. y = The amount of 25% salt water to be added to the tank. Then, x + y = 5 gallons Perfect. Now, equate the salt. How much salt will we get from 'x'? How much salt will we get from "y"? How much salt in total?

OpenStudy (anonymous):

x/10+y/4=(x+y)/5 4x+10y=8x+8y 4x=2y y=2x total of volume x+y=5 x+2x=5 3x=5,x=5/3,y=2x=10/3 I got this but my teacher told me I'm missing the complete final answer which I dont get.

OpenStudy (anonymous):

Nvm I see the problem. 10 percent 5/3 gallons 25 percent 10/3 gallons I forgot to put the percentage of each.

OpenStudy (anonymous):

thank you though !

OpenStudy (anonymous):

Can you help me with this one please.

OpenStudy (tkhunny):

You will have to show some proficiency in this process. You should start by making the scatter plot.

OpenStudy (anonymous):

I turn it in already I review it and redid it. Can you plx help me on this one? I attach the image. I already did the plotting for the points.

OpenStudy (tkhunny):

Already saw it and already commented on it. Make a scatter plot. No one can help you with that. You have to do it.